# Double Spending Analysis of Nakamoto Consensus for Time-Varying Mining Rates with Ruin Theory

### v1

@@ IV-A Reduction to Other Delay Models
Note that none of the [19, 22, 26, 25, 20] considers scaling the mining rate in order to make sure $\mathcal{T}$-block interval rule is satisfied. Finding the honest block generation rate in rigged model of [22] is too complex. On the other hand, a simple adjustment of $\lambda$ (resp. $\mu_{m}$) as $\frac{1}{(1-\beta)(\mathcal{T}-\Delta)}$ in [19] (resp. [26]) would make sure $\mathcal{T}$-block interval rule is satisfied. Similar adjustments can be made for [25]. However, these models also involve inspection paradox, hence, when we compare our results with these models, we simply take the corresponding ME-approximation with the reduction methods listed above. Thus, in the numerical examples, whenever we refer to the zero-delay or fixed-delay model, we obtain their related results using the reduction of $T$-matrix instead of using results directly from [22, 19] which do not originally obey $\mathcal{T}$-block interval rule. By using the $T$-matrix reduced version of their results, we make sure that $\mathcal{T}$-block interval rule is satisfied.
@@ V Numerical Results
We perform the following processing on raw data in several steps. Some nodes in the data seem to send inventory messages much later than usual and so we assume that they are light nodes that only echo messages and do not contribute to mining. Thus, in the first step, for a given parameter $\epsilon>0$, we find the $100(1-\epsilon)$-th percentile, or simply $\Delta(\epsilon)$ which is the cutoff delay that represents the largest of the block delays that occur $(1-\epsilon)$ of the time in the raw data, and subsequently eliminating all the block delays strictly larger than $\Delta(\epsilon)$. This step is needed to avoid overly skeptical estimates for the double spending probability.
@@
In Figure 3, we show how mining rate varies over time for both full data as well as congested network data, i.e., $\alpha(t)$, where we pick $\epsilon=0.01$, $K=27$ and $N^{\prime}\in\{8,32,128\}$ Notice that, especially in congested network situation, as delays are larger, eventually $T$-block interval rule results in easier puzzles. This in turn means at full honest power, i.e., when everyone receives the most recently mined block, $\alpha>1/600$. Eventually, this will also affect the double spending probability since $\beta$ is a constant fraction of full power $\alpha$ and it increases as well. In other words, during network congestion, as puzzle difficulty drops, adversary benefits from it.
(a) full data
@@
In Fig. 5, we compare the double spend probability $q$ obtained with the zero-delay model as well as the variable-delay model using the proposed ME distribution for three different values of the cutoff parameter $\epsilon\in\{0.01,0.001,0.0001\}$, where $\beta=0.2\alpha$ using the full range data. For the fixed-delay model, we plot both the $\Delta=10$ case as in [19], and additionally pick $\epsilon=0.01$ and find the corresponding $\Delta$ from full data ($\Delta\approx 300s$). Note that the choice of the cutoff parameter $\epsilon$ has a substantial impact on the double spend probabilities whose accurate estimation is key to the security. Further, the fixed-delay model either ends up almost the same as zero-delay model or gives extremely pessimistic values when only some of the nodes experience high delays.

### v2

@@ IV-A Reduction to Other Delay Models
Note that none of the existing works consider scaling the mining rate in order to make sure $\mathcal{T}$-block interval rule is satisfied. Finding the honest block generation rate in rigged model of [28] is too complex. On the other hand, a simple adjustment of $\lambda$ (resp. $\mu_{m}$) as $\frac{1}{(1-\beta)(\mathcal{T}-\Delta)}$ in [18] (resp. [24]) would make sure $\mathcal{T}$-block interval rule is satisfied. Similar adjustments can be made for [23]. However, these models also involve inspection paradox, hence, when we compare our results with these models, we simply take the corresponding ME-approximation with the reduction methods. Thus, in the numerical examples, whenever we refer to the zero-delay or fixed-delay model, we obtain their related results using the reduction of $T$-matrix instead of using results directly from [28, 18] which do not originally obey $\mathcal{T}$-block interval rule.
@@ V Numerical Results
We perform the following processing on raw data in several steps. Some nodes in the data seem to send inventory messages much later than usual. We assume, they are light nodes that only echo messages and do not contribute to mining. Thus, in the first step, for a given parameter $\epsilon>0$, we find the $100(1-\epsilon)$-th percentile, or simply $\Delta(\epsilon)$ which is the cutoff delay that represents the largest of the block delays that occur $(1-\epsilon)$ of the time in the raw data, and subsequently eliminating all the block delays strictly larger. This step is needed to avoid overly skeptical estimates for the double spending probability.
@@
In Figure 3, we show how mining rate varies over time for both full range data as well as congested network data, i.e., $\alpha(t)$, where we pick $\epsilon=0.01$, $K=27$ and $N^{\prime}\in\{8,32,128\}$. Notice that, especially in congested network situation, as delays are larger, eventually $T$-block interval rule results in easier puzzles. This in turn means at full honest power, i.e., when everyone receives the most recently mined block, $\alpha>1/600$. Eventually, this will also affect the double spending probability since $\beta$ is a constant fraction of full power $\alpha$ and it increases as well. In other words, during network congestion, as puzzle difficulty drops, adversary benefits from it.
(a) full range data
@@
In Fig. 5, we compare the double spend probability $q$ obtained with the zero-delay model as well as the variable-delay model using the proposed ME distribution for three different values of the cutoff parameter $\epsilon\in\{0.01,0.001,0.0001\}$, where $\beta=0.2\alpha$ using the full range data. For the fixed-delay model, we plot both the $\Delta=10$ case as in [18], and additionally pick $\epsilon=0.01$ and find the corresponding $\Delta$ from full range data ($\Delta\approx 300s$). Note that the choice of the cutoff parameter $\epsilon$ has a substantial impact on the double spend probabilities whose accurate estimation is key to the security. Further, the fixed-delay model either ends up almost the same as zero-delay model or gives extremely pessimistic values when only some of the nodes experience high delays.
Note that, the fixed delay model with $\Delta=10$ would correspond to a cutoff parameter of approximately $\epsilon=0.1$ in the delay data we use. The models that used $\Delta=10$ in the literature such as [19, Section III.C] acknowledge that higher delays are present in real world data but they can be seen as adversarial entities which is a pessimistic assumption. Therefore, it is not fair to compare a scenario of different adversarial powers with our model. Hence, we opted to present two cases of the fixed delay model: one where it is optimistic and ignores the remaining delays ($\epsilon=0.1$) and the other (more pessimistic) where the cutoff is $\epsilon=0.01$. Further, we opted to present Bitcoin results due to the abundance of reliable real world data and considerable interest from the research community and broader public. However, our model could be even more interesting for PoW blockchains that have smaller block interarrival times where the need of a better approximation of double spending probabilities is evident.
VI Discussion
In this paper, we introduced a ruin-theoretical analytical model of double spending using a matrix-analytical algorithm by which empirical network delays are used to construct a matrix-exponential distribution to model honest inter-mining durations, when the adversary hashrate is fixed, i.e., the adversary inter-mining periods are exponentially distributed with fixed rate. We proposed a novel method that incorporates the effect of arbitrary network delays on double spending probabilities, which was missing in the literature, even for fixed adversary hashrates. We leave the modeling of varying adversary hashrates for future research.
We used empirical network delays reported in the literature to construct a matrix-exponential distribution to model honest inter-mining durations. While doing so, in our numerical results, we made the assumption that all miners have the same hashrate, which may be a simplistic assumption. This assumption was attributed to the fact that we did not have access to data showing the hash powers associated with inventory messages. It is possible to re-run the model with a more detailed empirical network delay data that reports the hash powers associated with inventory messages.
Our model is heuristic, however, our goal is not to provide pessimistic bounds, but rather lay out a foundational framework, i.e., ruin theory, for more realistic estimation of double spending probabilities in PoW blockchains and more general scenarios. The accuracy depends on the assumption that future network delays follow the same pattern as observed in the past. However, the model can be re-run as new network delay data emerges. Taking into account of a finite collection of miner pools with varying hashrates and detailed peer-to-peer delays within the ruin-theoretical model is an interesting future research direction. We believe that the ruin-theoretical double spending model we introduced in this paper will lay out a foundation for more realistic and interesting blockchain scenarios, more general than that covered by the conventional gambler’s ruin framework. In fact, (8) can be modified using the iterated law of expectation to consider different pool hashrates and their respective delays to the other pools.
The ME-distribution model allows us to take a unifying view to many other problems studied extensively in the literature as sub-cases, such as zero-delay, fixed-delay, exponentially-fixed-delay, and more specifically variable-delays, the latter being the main scope of this paper. The order $m$ of the characterizing matrices of the ME distribution governs the complexity $\mathcal{O}(m^{3})$ of the approach since the algorithm involves the solution of matrix equations of size $m$ which is far larger than the parameter $k$ in the $k$-deep confirmation rule. In the model we propose for variable delays, $m=NK+1$ where $N$ is the discretization parameter and $K$ is the order of the CME distribution. In the numerical examples, we take $K=27$, $N$ up to 130, and the results converged within reasonable accuracy for these parameters, for which we have been able to obtain the double spending probabilities in seconds, with ordinary computing resources.